# Search-Based Software Engineering in the Landscape of AI Foundation Models

### v1

@@ 4.3 Search-based Debugging, Repair, and Evolution of FMs-Generated Artifacts
LLM-based fault localization and reproduction techniques generate fault reports, which are subsequently used for tasks such as debugging, program repair, and test evolution. Given that SBSE literature has well-established techniques for software maintenance and evolution, we argue that these techniques can be leveraged to enhance the process. For instance, search-based program repair techniques can be adapted to improve code as a next step after fault localization by LLMs. Moreover, once the code has been improved, another direction involves using SBSE to evolve tests based on the fault information.
5 Interplay between Search-based Software Engineering and Foundation Models
@@
A growing research trend is to explore the synergy between SBSE and LLMs, investigating how LLMs can enhance SBSE techniques and how SBSE can improve LLMs for SE tasks. One notable area is software development, where Gao et al. [15] introduced SBLLM, a framework for code optimization that leverages LLMs to apply evolutionary search techniques like crossover and mutation. For the application-specific use of search and LLMs, Shojaee et al. [37] presented LLM-SR, an approach that combines LLMs and evolutionary search to discover mathematical equations from the programs, demonstrating the potential to enhance the accuracy and efficiency with both LLMs and search techniques. Another prominent area of interest is exploring the potential of LLMs with SBST. In this direction, Dakhama et al. [12] introduced SearchGEM5, an approach that combines search-based fuzzing with LLMs to generate test cases for the Gem5 simulator. Later, Bruce et al. [6] presented an approach that utilizes LLMs to guide search-based fuzzing toward finding faults in ARM simulator software. Lemieux et al. [22] presented a technique called CodaMosa, which integrates a multi-objective search algorithm with a coding LLM to generate tests for Python code capable of achieving high coverage. Furthermore, Sapozhnikov et al. [33] presented TestSpark, a tool that utilizes the search-based capabilities of EvoSuite and LLMs’ textual capabilities to generate Junit tests. Subsequently, Abdullin et al. [1] conducted a comparative study of search-based tools, symbolic execution, and TestSpark configured with various LLMs. Their results demonstrated that while LLM-based tools underperformed in the case of code coverage, they outperformed in achieving higher mutation scores.
@@
While the synergistic potential between SBSE and FMs remains largely unexplored, early research reveals a promising interplay. However, it also highlights a significant performance limitation due to the substantial computational resources required by the integrated use of SBSE and FMs. In the following, we present a research roadmap and opportunities, exploring the potential of SBSE and FMs in advancing metaheuristics search in SBSE, overcoming FM-related challenges, solving SE problems, and addressing challenges in emerging domains.
5.1 Advancing Metaheuristics Search in SBSE
@@
As highlighted in the latest survey [42] and our literature analysis, FMs, specifically LLMs, can support various search algorithms in multiple ways. This includes using FMs to generate population, perform crossover and mutation, act as fitness functions, and support multi-objective optimizations. Given the advantages that FMs bring to the search process, one potential research direction is developing novel FM-inspired search algorithms. Furthermore, search-based algorithms, such as genetic algorithms (GAs), initiate the search process with a randomly generated population. During the search process, different selection strategies like tournament and rank-based selection are used to pick individuals from the current population to generate offspring for the next generation. For solving complex domain problems, intelligently selecting and reusing solutions individuals has demonstrated to be an efficient strategy [35, 36]. In this context, a promising research direction is using FMs to leverage domain knowledge for generating an optimized initial population, thereby enhancing the efficiency of the search process. For instance, for SBSE applied to emerging domains requiring mixed-mode capabilities, such as robotics and ADS, FMs like VLMs or MMs can support generating the initial population. This could significantly broaden SBSE’s applicability to diverse, complex domains with cutting-edge challenges. Moreover, during the search, FMs can be used to calculate fitness, depending on the SBSE problem at hand. For example, in ADS testing, FMs can be queried to assess whether a scenario is realistic. This information can then be used to define fitness functions that guide search algorithms for SBSE problems. Thus, more research is needed to investigate which SBSE problems can utilize FMs to calculate fitness during the search. Another significant challenge is that search algorithms often generate invalid solutions that require repairs. To address this, FMs can repair solutions during the search, ensuring the search is properly guided. To this end, new research methods based on FMs are needed to prompt FMs to repair solutions when issues are encountered during the search.
5.2 Addressing FM Challenges
A key challenge of FMs is non-determinism, where the same input can produce varying outputs, affecting consistency and reliability. This non-determinism leads to several challenges in applying FMs in SE development phases. For instance, in CI/CD and DevOps, variability in generated configurations may lead to inconsistencies in software releases. SBST can complement handling non-determinism in FMs by search-based robustness testing techniques with the objective of exploring inputs leading to high model variability. Another direction in which SBSE can be applied is to find the best configurations for software and hardware resources that minimize variability in model behavior. Another challenge of FMs is their inherent uncertainty. Since FMs are trained on vast amounts of data, fine-tuning them also requires a significant amount of data, and the quality of such data (e.g., redundant features or imbalanced data) can be a potential source of uncertainty. Such uncertainties can significantly affect model interference, leading to ineffective performance for several SE phases. SBSE can be applied to select an optimal set of features, which will result in effectively managing data-related uncertainty and efficient training/fine-tuning of FMs. Furthermore, a key challenge in FMs is hallucination, where the model generates entirely fabricated or factually incorrect information [32], posing a significant concern in FM applications. As a result, detecting and mitigating hallucinations has become a key focus of ongoing research [8]. In this regard, we foresee significant potential for SBSE to effectively address the challenge of hallucinations in FMs. Specifically, SBSE can complement black-box methods for hallucination detection, as highlighted by the need for such methods [8]. By leveraging a search-based approach, a search algorithm can iteratively generate a diverse population of prompts, collect corresponding responses from FMs, and use a fitness function to continuously assess response variations, enabling more precise hallucination detection.
5.3 Solving SE Problems
@@
SBSE has been applied to many software requirement engineering (RE) problems, including requirement selection and prioritization, requirement satisfaction, and fairness analysis in requirements. Software requirements are typically elicited and documented in a textual format. FMs, specifically LLMs with text generation, comprehension, and analysis capabilities, can complement search-based RE in many ways. For instance, in the case of requirements selection, LLMs can be provided with software requirements and then act as a fitness function during search optimization for solving the next release problem. Moreover, for requirements prioritization with multi-objective search, LLMs can be used to analyze conflicting objectives and select Pareto-dominant solutions. In addition to RE, SBSE techniques can be applied to test FMs used to solve SE problems (e.g., code generation and test generation). Taking the example of FMs generating source code, the aim is to test whether FMs generate correct code. For this purpose, SBST can be applied to generate test cases that prompt FMs to generate source code. The generated source code can then be used to compute a fitness function, guiding the search to identify faults that make FMs generate the wrong code. To this end, with carefully designed fitness functions, search-based testing techniques can be used to identify faults in FMs for various software engineering tasks. Once faults are identified using SBSE approaches, they can also be employed to repair FMs. This can be achieved by defining a fitness function to generate patches (e.g., fine-tuning parameters) for FMs until the faults are resolved. In software testing, graphical user interface (GUI) testing presents a compelling application for integrating SBSE and FMs. Modern software applications are designed with advanced GUIs to facilitate user-friendly interactions. GUI testing is essential to ensure functionality, visual consistency, compatibility, and usability. The available tools for GUI testing, like GUITAR [30] and Selenium [7] require test scripts to execute tests. Therefore, combining SBST with MMs, which possess mixed-mode capabilities, such as textual and graphical elements, presents an intriguing direction for future research.
5.4 FM-Integrated SBSE for Emerging Domains
FMs integrated with SBSE can be used to solve complex and dynamic nature problems in emerging domains. For example, in system-level testing of live Internet of Things (IoT) applications, the comprehension and analytical capabilities of FMs (such as LLMs or MMs), combined with SBST, can enable testers to perform online testing in an organized and intelligent manner. Similarly, VLMs and MMs can facilitate the safety analysis of self-adaptive robotics by processing mixed-mode data from cameras, sensors, and actuators. Overall, we envision that integrating FMs with SBSE will offer substantial support across the software engineering lifecycle in diverse domains, including cyber-physical systems, smart grids, and autonomous vehicles.
6 Related Works

### v2

@@ 7.2.5 Search-based Debugging, Repair, and Evolution of FMs-Generated Artifacts
LLM-based fault localization and reproduction techniques generate fault reports, which are subsequently used for tasks such as debugging, program repair, and test evolution. Given that the SBSE literature has well-established techniques for software maintenance and evolution, we argue that these techniques can be leveraged to enhance the process. For instance, search-based program repair techniques can be adapted to improve code as a next step after fault localization by LLMs. Moreover, once the code has been improved, another direction involves using SBSE to evolve tests based on the fault information.
8 Interplay between SBSE and FMs
In this section, we explore the interplay between SBSE and FMs, highlighting their synergistic potential. We begin with an overview of the current state of the art and then discuss the key challenges and opportunities for further research.
8.1 Current Landscape
An increasing research trend is to explore the synergy between SBSE and LLMs, investigating how LLMs can enhance SBSE techniques and how SBSE can improve LLMs for SE tasks. One notable focus is software development, where Gao et al. [56] introduced SBLLM, a framework for code optimization that leverages LLMs to apply evolutionary search techniques like crossover and mutation. Similarly, Bouras et al. [20] presented LLM-based crossover operators for genetic improvement of the source code. Moreover, Wang et al. [181] proposed a masking mutation operator for genetic improvement that uses LLM code completion abilities for context-aware code changes during the mutation process. Another area is software genetic improvement (GI), where Kang and Yoo [77] presented the initial concept of generating mutants from LLMs to optimize the process of GI. Similar to this work, Brownlee et al. [22] employed LLMs as mutation operators to increase diversity in the search process for GI. In the case of automated program repair, Bin Murtaza et al. [18] presented SemiAutoFL, a tool that combines genetic programming with LLMs to improve fault localization and bug fixing. For the application-specific use of search and LLMs, Shojaee et al. [160] presented LLM-SR, an approach that combines LLMs and evolutionary search to discover mathematical equations from the programs, demonstrating the potential to enhance the accuracy and efficiency with both LLMs and search techniques.
@@ 8.1 Current Landscape
Another prominent area of interest is exploring the potential of LLMs with SBST. In this direction, Dakhama et al. [36] introduced SearchGEM5, an approach that combines search-based fuzzing with LLMs to generate test cases for the Gem5 simulator. Building on this work, Dakhama et al. [37] introduced SearchSYS, an automated approach and tool designed to generate diverse test inputs using LLMs and code mutants through search-based fuzzing. Later, Even-Mendoza et al. [45] presented an approach that utilizes LLMs to guide search-based fuzzing toward finding faults in ARM simulator software. Lemieux et al. [89] presented a technique called CodaMosa, which integrates a multi-objective search algorithm with a coding LLM to generate tests for Python code capable of achieving high coverage. Furthermore, Sapozhnikov et al. [146] presented TestSpark, a tool that utilizes the search-based capabilities of EvoSuite and LLMs’ textual capabilities to generate JUnit tests. Subsequently, Abdullin et al. [1] conducted a comparative study of search-based tools, symbolic execution, and TestSpark configured with various LLMs. Their results demonstrated that while LLM-based tools underperformed in the case of code coverage, they outperformed in achieving higher mutation scores. In the ADS domain, Tang et al. [171] used LLMs to derive logical scenarios from accident reports and then applied multiobjective search to generate critical test scenarios.
8.2 Roadmap
@@
While the synergistic potential between SBSE and FMs remains largely unexplored, early research reveals a promising interplay. However, it also highlights a significant performance limitation due to the substantial computational resources required by the integrated use of SBSE and FMs. In the following, we present a research roadmap and opportunities, exploring the potential of SBSE and FMs to advance metaheuristics search in SBSE and solve challenges in SE and emerging domains.
8.2.1 Advancing Metaheuristics Search in SBSE
As highlighted in the latest survey [193] and our literature analysis, FMs, specifically LLMs, can support various search algorithms in multiple ways. This includes using FMs to support single and multi-objective optimizations, act as fitness functions, and perform crossover and mutation.
Advancing search algorithms.
@@ 8.2.1 Advancing Metaheuristics Search in SBSE
Given the advantages that FMs bring to the search process, one potential research direction is developing novel FM-inspired search algorithms. Furthermore, search-based algorithms, such as genetic algorithms, initiate the search process with a randomly generated population. During the search process, different selection strategies like tournament and rank-based selection are used to pick individuals from the current population to generate offspring for the next generation. For solving complex domain problems, intelligently selecting and reusing solutions individuals has demonstrated to be an efficient strategy [148, 155]. In this context, a promising research direction is using FMs to leverage domain knowledge for generating an optimized initial population, thereby enhancing the efficiency of the search process. For instance, for SBSE applied to emerging domains requiring mixed-mode capabilities, such as robotics and ADS, FMs like VLMs or MMs can support generating the initial population. This could significantly broaden SBSE’s applicability to diverse, complex domains with cutting-edge challenges.
Advancing fitness functions.
During the search, FMs can be used to calculate fitness, depending on the SBSE problem at hand. For example, in ADS testing, FMs can be queried to assess whether a scenario is realistic [192]. This information can then be used to define fitness functions that guide search algorithms for SBSE problems. Thus, more research is needed to investigate which SBSE problems can utilize FMs to calculate fitness during the search. Another significant challenge is that search algorithms often generate invalid solutions that require repairs. To address this, FMs can repair solutions during the search, ensuring the search is properly guided. To this end, new research methods based on FMs are needed to prompt FMs to repair solutions when issues are encountered during the search.
Advancing search operators.
In a recent study, Wu et al. [193] categorized various LLM-based search operators, such as crossover and mutation, for single and multi-objective algorithms. However, the application of such operators in SBSE remains undetermined. Therefore, one potential research direction is to explore LLM-based search operators in the SBSE context. Another research direction is to determine search operators at runtime using FMs, like VLMs and MMs, and apply SBSE to solve problems for dynamic systems. For example, when dealing with ADS data in image format, VLMs can be used to perform crossover and mutation for the generation of image-based test scenarios [57]. Similarly, for ADS data in audio or video formats, search operators based on SAMs and MMs can enhance test scenarios considering the mixed-mode nature of data.
8.2.2 Solving SE Problems
SBSE has been applied to many software requirement engineering (RE) problems, including requirement selection and prioritization, requirement satisfaction, and fairness analysis in requirements. Software requirements are typically elicited and documented in a textual format. FMs, specifically LLMs with text generation, comprehension, and analysis capabilities, can complement search-based RE in many ways. For instance, in the case of requirements selection, LLMs can be provided with software requirements and then act as a fitness function during search optimization for solving the next release problem. Moreover, for requirements prioritization with multi-objective search, LLMs can be used to analyze conflicting objectives and select Pareto-dominant solutions.
@@ 8.2.2 Solving SE Problems
In addition to RE, SBSE techniques can be applied to test FMs used to solve SE problems (e.g., code generation and test generation). Taking the example of FMs generating source code, the aim is to test whether FMs generate correct code. For this purpose, SBST can be applied to generate test cases that prompt FMs to generate source code. The generated source code can then be used to compute a fitness function, guiding the search to identify faults that make FMs generate the wrong code. To this end, with carefully designed fitness functions, search-based testing techniques can be used to identify faults in FMs for various SE tasks. Once faults are identified using SBSE approaches, they can also be employed to repair FMs. This can be achieved by defining a fitness function to generate patches (e.g., fine-tuning parameters) for FMs until the faults are resolved.
In software testing, graphical user interface (GUI) testing presents a compelling application for integrating SBSE and FMs. Modern software applications are designed with advanced GUIs to facilitate user-friendly interactions. GUI testing is essential to ensure functionality, visual consistency, compatibility, and usability. The available tools for GUI testing, like GUITAR [122] and Selenium [23] require test scripts to execute tests. Therefore, combining SBST with MMs, which possess mixed-mode capabilities, such as textual and graphical elements, presents an intriguing direction for future research.
9 Challenges of Empirical Evaluations
When proposing a novel technique to address an SE problem, typically experiments are carried out on a set of artifacts, with comparisons made with the current state-of-the-art. If better results are obtained, then a novel contribution to the current body of knowledge has been made. However, to ensure that a novel technique $X$ is indeed better than the current state-of-the-art $Y$, experiments need to be fair.
Typically, in SBSE experiments, and in experiments with search algorithms in general, compared techniques are run with the same resources (e.g., 1 CPU-core) for the same search budget. This could be a timeout (e.g., 2 minutes or 1 hour), or the same number of evaluated solutions (recall Section 3.1). In other words, algorithms and techniques/tools are compared by trying to use the same amount of computational resources. If you do not have fair comparisons, you cannot really claim that a novel technique is better than the current state-of-the-art.
When using FMs or combining FMs with SBSE techniques, the same requirements still hold: experiments must be fair. However, how to ensure that the same amount of resources is employed in the case of FMs? If you use a commercial FM, typically, you will be making an API call over the internet by sending a prompt as a payload. This payload would then trigger computations on a remote server, usually running on high-end GPUs. Then the response would be sent back. Such round-trip communication toward a remote server takes a non-zero amount of time. Furthermore, what are the specs of those high-end GPUs? How much electricity did they consume to answer the prompt? How do those values compare with the specs of a consumer-grade laptop where you can run SBSE techniques?
If a remote, commercial FM requires $N$-times the electricity and hardware capabilities compared to an SBSE technique, a more fair comparison would be to run the latter in parallel with $N$ CPU-cores. If a laptop/server does not have enough CPUs, then, as commercial FMs are run on remote servers, you could rent cloud resources (e.g., from Amazon Web Services or Azure) to run the SBSE technique in parallel. If you can pay money for a subscription to use a commercial FM (e.g., ChatGPT), then you can use the same amount of money to rent cloud resources to run an SBSE technique in parallel. Still, how to reliably determine a fair value for $N$?
The use of remote, commercial FMs is problematic in this context. An alternative would be to use an open-source FM that can be run locally (e.g., via ollama [^1] https://ollama.com/), as it makes evaluating its computational costs more feasible. However, without the proper hardware (e.g., GPUs) results could be quite poor compared to commercial, remotely run FMs. The need for specialized, more expensive hardware (compared to a consumer-grade laptop) is not necessarily a showstopper. Such costs could be justified based on the improvements obtained by solving the SE problems at hand. Still, how to make proper fair comparisons when different techniques might require different kinds of hardware, e.g., GPUs vs. CPUs? There is no easy answer here, apart from possibly checking their power consumption, which is not necessarily trivial to infer reliably.
However, the use of a local FM solves another critical issue: replicability. A remote, commercial FM could change at any time, like any other service provided on the internet. Given an experiment involving such a remote FM, it might not be possible to replicate such a study if the FM has changed meanwhile [144]. Lack of replicability is a major issue in scientific research [127, 34]. A result that cannot be replicated has limited scientific value. On the other hand, if the FM is an open-source one that can be run locally, experiments can be attempted to replicate, as long as similar hardware is employed. This does not mean that commercial FMs should not be used in scientific research, especially if they are the ones that give the best results. Still, for sound results, experiments should be run as well with a local FM, which can then be used as a point of reference in replicated studies and comparisons.
Note that the concept of “replicability” is different from the concept of “non-determinism”. If you prompt an FM twice, you might get different results even if the prompt is exactly the same. But this is exactly the same issue as for SBSE techniques, which most of the time are based on randomized algorithms. In those cases, experiments need to be repeated several times (e.g., 10 or 30 times), and the appropriate statistical tests should be used to analyze the results [9].
How to properly compare, in a fair manner, hybrid techniques involving both FMs and SBSE techniques is not going to be trivial. Detailed guidelines will need to be established in the SBSE/FM research community to address this important research problem that impacts all empirical studies on this subject.
10 SBSE and FMs: Research Horizon for 2030
In this section, we explore the future of SBSE in the context of FMs as a disruptive technology. We then highlight the key challenges and emerging opportunities that are likely to form the research horizon for 2030.
10.1 The Future of SBSE
We analyze the impact of FMs on SBSE through McLuhan’s Tetrad [111], a framework designed to examine the disruptive effects of emerging technologies. The tetrad consists of four key aspects: what a new technology enhances, retrieves, renders obsolete, and reverses when pushed to extremes. In the context of SBSE, we follow the disruptive research playbook by Storey et al. [166] and apply McLuhan’s Tetrad to assess the transformative role of FMs. Specifically, we explore four questions: (i) How do FMs enhance SBSE? (ii) What do FMs retrieve that were previously overlooked in SBSE? (iii) Which traditional SBSE techniques and tools may become obsolete due to FMs? And (iv) what in SBSE reverses due to over-reliance on FMs? The outcomes of this analysis are summarized in Figure 3 and elaborated in the discussion that follows.
Figure 3: An overview of the tetrad illustrating the disruptive effects of FMs on SBSE.
It is essential to note that, although this analysis is based on our review of the state-of-the-art, it does not provide definitive predictions of the future.
10.1.1 Enhances
FMs can significantly enhance automation in SBSE by reducing manual efforts across various design elements, such as population generation, solution encoding, fitness function design, crossover and mutation operators, and selection mechanisms. This will drive the development of novel search techniques, the creation of novel FM-inspired search algorithms, and improvements over existing methodologies, such as genetic programming, by providing advanced reasoning and contextual understanding. The multimodal capabilities of FMs will augment conventional problem-solving in SBSE, enabling the integration of diverse data types (e.g., text, images, audio, and video) and facilitating broader applicability to emerging domains such as ADS and robotics. These capabilities will also allow domain experts to participate in the SBSE process without requiring in-depth knowledge of SBSE, fostering interdisciplinary collaboration in diverse fields, including healthcare, finance, and transportation. Furthermore, the development of novel search techniques powered by FMs will enhance existing tools and frameworks, such as jMetal [43], by integrating FM-based features like automated prompt generation, multimodal data handling, and intelligent solution evaluation. As a result, these tools will facilitate decision-making in SBSE by providing advanced analytics and insights, such as identifying trade-offs in multi-objective optimization problems or suggesting alternative solutions, as well as customizing SBSE solutions to specific domain requirements.
10.1.2 Retrieves
FMs will bring back human-centric interaction to SBSE by enabling intuitive problem-solving through natural language interfaces. This shift allows users, even those without technical expertise, to define search problems and interpret solutions more easily. By reintroducing accessible, conversational workflows, FMs will make SBSE more inclusive, recalling earlier human-focused problem-solving methods that prioritized simplicity and user engagement. Another interesting perspective is that FMs will revive the creative and exploratory nature of problem-solving by generating novel solutions, retrieving overlooked alternatives, and providing innovative suggestions. Through contextual understanding, FMs will enable solutions that are based on domain-specific knowledge, reintroducing the principles of domain-driven design with enhanced scalability and intelligence. Moreover, FMs, such as VLMs, will reintroduce the importance of graphical and visual representations in SBSE. They will enhance both problem representation for various SE design artifacts, such as UML diagrams and flowcharts, and solution representation, such as visually illustrating Pareto fronts for multi-objective optimization. Lastly, FMs will utilize large repositories of historical data to retrieve valuable knowledge, identifying patterns in legacy code and historical project data. This capability will play an important role in reviving legacy systems by facilitating tasks such as understanding older codebases and enabling efficient system modernization.
10.1.3 Reverses
When pushed to extremes, FMs have the potential to overshadow SBSE, thus losing the significance of SBSE. This will shift the key focus from SBSE problem-solving mechanisms to prompt engineering as the primary means of achieving outcomes. As a result, the significance of SBSE, as the creative and systematic design of solutions, may be replaced by reliance on pre-trained language models. Over-reliance on FMs will also transfer their inherent limitations and risks, such as issues of non-determinism, hallucinations, and ethical concerns, to every SBSE application. Furthermore, the computational demands associated with using FMs, especially for large-scale problems, may lead to resource constraints, leading to sustainability concerns for real-world SBSE applications. An increasing dependence on FMs could also suppress innovation in SBSE by averting the development of novel algorithms and techniques, as the focus shifts to applying FMs rather than advancing SBSE with new methods. Over time, this could result in a loss of SBSE expertise among software engineers, as basic knowledge and skills are no longer exercised in practice.
10.1.4 Obsolesces
The integration of FMs into SBSE will render many manual design practices obsolete, such as solution encoding and fitness functions, as these tasks will become increasingly automated. Traditional methods that rely on predefined or fixed search spaces will also become redundant, as FMs will enable intelligent and adaptive search space definitions based on context understanding. Alongside these changes, several search algorithms and heuristics designed for simpler problems may become obsolete or lose relevance. For example, heuristics for model-based test data generation using constraint specifications (such as [3, 155]) will no longer be required, as FMs, particularly LLMs, can directly generate test data using these specifications. Similarly, traditional search algorithms like the Alternating Variable Method (AVM), which were commonly used for specific SBSE problems, may be reshaped or rendered obsolete entirely. The obsolescence of such algorithms and heuristics will, in turn, affect the tools and frameworks that rely on them, which will need to adapt to FM-introduced changes or risk becoming irrelevant. Furthermore, the need for domain-specific tailoring of search-based solutions may become obsolete, as FMs provide generalization capabilities with the flexibility to adapt to a diverse range of problems and domains.
10.2 Challenges and Opportunities Ahead
This section begins by summarizing the key challenges and then presents potential opportunities in emerging domains.
10.2.1 Overview of Core Challenges
LABEL:tab:keychallenges provides a detailed summary of the key challenges discussed in Sections 4, 5, 7, 6, 8 and 9. It highlights open research avenues that must be addressed to advance SBSE in the context of FMs, particularly in solving complex multidisciplinary problems.
Table 1: Key challenges in achieving a synergistic potential between SBSE and FMs, along with challenges in conducting empirical evaluations involving both.
FMs for SBSE Design
✦ Automating the design and implementation of fitness functions for SBSE problems.
✦ Designing search operators for complex domain problems involving multimodal data.
✦ Automating solution encoding for the varying nature of problems.
✦ Defining appropriate search spaces for SBSE to balance efficiency and solution quality.
✦ Generating partial or complete SBSE implementations to enhance automation and efficiency.
✦ Automating the testing, debugging, and repair of SBSE implementations.
FMs for SE Artifacts Generated with SBSE
✦ Interpreting and transforming SBSE-generated artifacts into textual or graphical formats.
✦ Static analysis of SBSE-generated artifacts for bug detection, test effectiveness, and program verification.
✦ Supporting Pareto analysis and pruning for multi-objective SE problems, particularly for high-dimensional trade-offs and intuitive decision-making.
✦ Validating and improving the quality of SBSE-generated artifacts, such as test scenarios, models, and code.
✦ Context-aware input generation, flaky test detection, and repair in delta debugging.
SBSE for FMs
✦ Managing non-determinism in FMs to ensure consistent and reliable outputs.
✦ Effectively handling data-related uncertainties in FMs to improve model performance.
✦ Detecting and mitigating hallucinations in FMs.
✦ Optimizing the computationally expensive fine-tuning process of FMs, particularly for evolving SE requirements.
✦ Balancing trade-offs in automated prompt optimization.
SBSE for SE Artifacts Generated with FMs
✦ Refining FM-generated code to address compiler errors, resolve conflicts, and integrate into codebases.
✦ Improving test quality, coverage, and diversity for FM-generated tests, while addressing flakiness and optimizing regression testing.
✦ Enhancing FM-generated models (e.g., UML diagrams) for quality, correctness, and constraint validation.
✦ Effectively optimizing FM-generated software configurations for functional and non-functional characteristics.
✦ Debugging and repair of FM-generated artifacts.
✦ Dealing with continuous evolution of FMs.
Interplay between SBSE and FMs
✦ Developing FM-inspired search algorithms to enhance initialization, fitness calculation, and solution repair during the search process.
✦ Using FMs to support SBSE in traditional SE problems, such as requirement prioritization, fault identification, and repair for SE tasks, including GUI testing.
✦ Effectively integrating FMs with SBSE to tackle complex problems in emerging domains.
Empirical Evaluations involving both SBSE and FMs
✦ Ensuring fair comparisons between FM-based and SBSE techniques, considering differences in computational resources and energy consumption.
✦ Determining appropriate scaling factors for resource allocation to achieve fairness in experiments.
✦ Addressing the replicability of experiments with remote, commercial FMs, which can change unpredictably over time.
✦ Balancing the use of commercial FMs for best results with local, open-source LLMs for replicability in scientific research.
✦ Managing non-determinism in FMs and SBSE techniques through repeated experiments and proper statistical analysis.
✦ Establishing detailed guidelines for fair and reliable comparisons of hybrid techniques involving both FMs and SBSE.
10.2.2 Opportunities in Emerging Domains
FMs integrated with SBSE can be used to solve complex and dynamic nature problems. In the following, we present opportunities to take advantage of the synergistic potential between SBSE and FMs in some emerging domains.
Autonomous Driving Systems.
SBSE has been extensively applied to Autonomous Driving Systems (ADS), addressing various aspects of software engineering, from requirements to testing [90, 103, 173, 88, 170, 25, 55, 5, 129]. Since ADS rely on mixed-mode data from various sensors such as LIDAR, radar, and cameras, search-based techniques have been integrated with deep learning methods to process these datasets, particularly in simulation environments [15]. However, ADS operate in highly dynamic environments having uncertain conditions such as varying road surfaces, unpredictable weather, and fluctuating traffic patterns, which creates a significant reality gap between simulations and real-world scenarios [192, 170]. Therefore, considering LLMs’ analytical capabilities, recent efforts have begun exploring LLMs to assist in bridging the simulation-to-reality gap [191]. Beyond LLMs, other FMs, such as VLMs and MMs, due to their context understanding and reasoning abilities, offer promising opportunities to enhance SBSE techniques in the ADS domain. One potential direction is to explore the integration of SBSE-based requirement selection and prioritization with MMs for requirement validation and traceability, enabling seamless processing of ADS requirements and multimodal sensor fusion data. Another direction is utilizing MMs to generate code directly from ADS requirements and design artifacts, followed by employing SBSE-driven automated program improvement techniques to optimize the generated code. In software testing, integrating FMs with SBSE presents opportunities such as generating realistic test scenarios for ADS, optimizing regression testing for evolving software, detecting and localizing bugs, and enhancing security testing to handle arising threats and vulnerabilities.
Robotics.
Autonomous robots employ adaptation mechanisms such as the MAPE-K loop and AI-based data analysis techniques [81] to autonomously monitor, analyze, and adapt to unpredictable dynamic environments and unforeseen events while interacting with humans and surrounding objects [87]. These robots can be categorized into various types depending on the application domain, such as industrial robots for automated manufacturing, service robots for office and home assistance, and healthcare robots for medical assistance. Similar to ADS, these robots also process multimodal sensor and actuator data, leading to similar opportunities for integrating SBSE with FMs.
A recent shift in robotics has seen the development of robotic foundation models, including fine-tuned LLMs and VLMs for robot-specific tasks and vision action language (VLA) models designed for robots [180, 179, 39]. These embodied robotic foundation models enable robots to operate in task-specific ways within closed environments while collaborating closely with humans. However, such interactions introduce several uncertain situations, including unpredictable human behavior, environmental changes, and task variability, raising significant reliability concerns [154, 47]. In this context, SBST presents promising opportunities for testing the robustness and reliability of robotic foundation models. A potential direction is to use FMs, such as VLMs and VLAs, to assist SBST in generating and simulating unpredictable human behavior, such as sudden gestures or commands, to evaluate the robot’s ability to adapt and respond safely. Due to the black-box nature of robotic foundation models, diagnosing faults in these models is challenging [61]. Therefore, another direction is to integrate FMs with SBST to detect and localize faults in robotic behaviors by analyzing multimodal data, such as correlating error logs with specific robot tasks or actions to identify inconsistencies. Furthermore, given the computational and time constraints associated with running extensive tests on robotic foundation models, a promising area of research lies in developing search-based test selection and prioritization techniques enhanced by multimodal FMs to focus on the most critical and impactful test cases.
AI-enabled Systems.
Search-based approaches have been extensively used for testing [78, 203, 140, 205, 141] and repair [162, 96, 167, 174, 95] of ML components like deep neural networks (DNNs) used in AI-enabled systems like ADS and robotic systems, for performing tasks like perception. In test generation, a typical approach is to generate adversarial inputs; however, such inputs may be unrealistic and do not occur in practice in the real world. To tackle this issue, for image classifiers, VLMs could help in generating more realistic inputs by analyzing the scene and adding and/or modifying elements that could lead to misclassification. DNN repair is a popular search-based approach that modifies a few network parameters (e.g., weights) with the goal of fixing some wrong behaviors by not degrading previously correct behaviors. The repair needs to be preceded by a fault localization phase [162, 44, 42, 107, 128] that identifies the suspicious components of the network that should be repaired. While for some DNN models, like classifiers, fault localization can be quite accurate, for other models, like regression models [156] or DNN controllers [106, 105, 104], the fault localization is quite challenging. In this case, VLM models could be used to analyze the scene and LLMs to reasons about possible causes of the violations; for example, a collision of an AI-enabled system guided by a DNN controller, could be analyzed by a VLA to identify critical frames; then, an LLM, by considering the system specification, could try to understand at which time point the controller gave a wrong control decision.
Internet of Things.
Internet of Things (IoT) applications consist of heterogeneous devices, often geographically distributed, along with third-party services, all interconnected through cloud-based architectures [150]. System-level testing of such applications with physical devices in the loop is often infeasible, making digital twins a preferred alternative for simulating these devices during testing [151, 152]. However, search-based testing of live IoT applications, especially after each upgrade during CI/CD practices, introduces several significant challenges [149, 153]. First, running a large number of test cases after every release during CI/CD can lead to service blocking from the servers of third-party device and service providers, potentially disrupting critical services. Second, fault localization becomes particularly challenging due to the interconnected nature of IoT systems. It is difficult to determine whether a specific failure originates from the application under test, connected devices, or third-party services. Third, bug reproduction is also challenging, especially for failures that occur irregularly or depend on specific runtime conditions that are difficult to replicate in testing environments. Fourth, identifying regressions, where previously resolved bugs reappear or new issues arise due to updates, is another major challenge, particularly in dynamic and evolving IoT ecosystems where changes in one component can have cascading effects across the system. To address these challenges, more research is needed to explore the potential of integrating FMs with SBST for IoT applications. FMs, with their advanced analytical capabilities and contextual understanding, can enable testers to perform online system testing in a largely automated, systematized, and intelligent manner. For example, FMs can analyze complex interdependencies between IoT devices and third-party services to improve fault localization by identifying the root cause of failures with greater accuracy. In addition, FMs can enhance search-based test selection and prioritization during CI/CD by analyzing mixed-mode data from IoT devices, their digital twins, and interconnected third-party services. For bug reproduction and regression identification, developing FM-based interactive tools to assist testers in recreating failure scenarios and identifying root causes presents a promising avenue for research. Furthermore, given the maturity and extensive study of search-based techniques, integrating them with recent advancements in applying LLMs for testing purposes offers another exciting direction for future exploration.
Quantum Computing and SBSE.
SBSE has also been applied to more advanced computing topics, such as quantum computing. For instance, recent works have demonstrated the use of SBSE to address quantum SE problems, including test generation [182, 183], mutation testing [184], and learning noise models for quantum computers [117]. However, such applications remain limited, and there are ample opportunities to apply SBSE to other phases of quantum SE, such as software modeling, code generation, and test optimization.
Moreover, current SBSE approaches primarily run on classical computers, although there is an increasing body of work exploring how SBSE can also be performed on quantum computers. To this end, the field of quantum search-based SE is emerging, where quantum search and optimization algorithms are applied to solve problems in classical software engineering [185]. However, such work remains minimal, and further research is needed. Along the same lines, the application of quantum search and optimization to solve quantum SE problems (i.e., quantum search-based quantum SE) remains an almost untouched area of research. We anticipate that by 2030, more research at the intersection of quantum computing and SBSE will emerge, presenting numerous opportunities for researchers.
11 Related Work